OpenStudy (anonymous):

Solve 3(2x - 8) = 12 You just Distribute, right?

OpenStudy (anonymous):

Distribute, then solve for x

OpenStudy (anonymous):

you can divide by 3 as a first step is you like, and get \(2x-8=4\)

OpenStudy (anonymous):

no need to multiply out since 3 is a factor of both sides

OpenStudy (anonymous):

of course you will get the same answer if you distribute first \[2x-8=4\] \[2x=12\] \[x=6\] vs \[6x-24=12\] \[6x=36\] \[x=6\] your choice

OpenStudy (anonymous):

That is what I got, when I worked it out! I was right! Thanks for checking and helping! :)